OpenStudy (anonymous):

Check my answer Solve the following system: y = x + 3 2x + y = 9

OpenStudy (anonymous):

(2, 5) <-- (5, 2) (-2, 5) (2, -5)

OpenStudy (shinalcantara):

y = x + 3 ----equation 1 2x + y = 9 ----equation 2 plugging in equation 1 to 2 2x + x + 3 = 9 3x = 9 - 3 3x = 6 x = 2 substitute to equation 1 y = x + 3 y = 2 + 3 y = 5

OpenStudy (anonymous):

thank you :)